OpenStudy (anonymous):

1+cosx= what identity?

OpenStudy (jango_in_dtown):

\[1+\cos 2x=2 \cos ^{2}x\]

OpenStudy (jango_in_dtown):

now instead of x pur x/2 and you will get the identity

OpenStudy (anonymous):

well the problem is \[\frac{sinx }{ sinx }+\frac{ sinx }{ 1+cosx }=2cscx\]

OpenStudy (jango_in_dtown):

you need to prove just??

OpenStudy (jango_in_dtown):

what us cscx? cosec x?

OpenStudy (jango_in_dtown):

@kirarittgers

OpenStudy (anonymous):

No it means 1/sinx

OpenStudy (jango_in_dtown):

please post the problem once again

OpenStudy (anonymous):

Well I'm on my phone now so i can't do an equation let me write it out then upload the picture